OpenStudy (anonymous):

solve for y 2(3x-1)+7y=5y-(x+3)

OpenStudy (anonymous):

2(3x-1)+7y=5y-(x+3) 7y-5y=-(x+3)-2(3x-1) simplify it

OpenStudy (anonymous):

just expand and make y the subject. so for the first one, you have: 6x-2+7y=5y-x-3 7x+1=-2y y= (7x+1)/-2 it's the same method for the second problem. :)